IIT JEE Entrance Exam

A block of mass 4 kg is placed on another block of mass 5kg and block B rests on a smooth horizontal surface. If the maximum force that can be applied of block A so that both the blocks move together is 12N, the maximum force that can be applied on block B for the blocks to move together is?
Please explain properly.

Kavya Singh
7 Years agoGrade

3 Answers

Arun
7 Years ago
When the blocks move together the acceleration will be same.  block A = mass4 kg and If f is the force of friction between A and B
F- f = mb-a  
And f = maa...  
So a = f/ma = 12/4 = 3m/s2
so F-12 = 8*3 = 24
F = 12+24 = 36 N
Kavya Singh
7 Years ago
Answer is given to be 27 newton as in the archive of jee mains from fiitjee.
 This question is asked in jee main 2014.
Arun
7 Years ago
Dear Kavya
 
I have not got the question in jee mains 2014 paper.
 
Please check and I don't think that 27 N is correct
 
Regards
Arun